OpenStudy (anonymous):

t+u=9 10t+u+u=24, solve to fins t and u

OpenStudy (asnaseer):

:/ if you really want to learn here you should take the advise I gave you in the previous question: http://openstudy.com/updates/4fe2457be4b06e92b87139bf

OpenStudy (anonymous):

yeah i saw the video

OpenStudy (asnaseer):

there is more than one video there...

OpenStudy (asnaseer):

that site also has a section with sample questions to practice on.

OpenStudy (anonymous):

oh really did't know that

OpenStudy (anonymous):

thing is that i need to turn in the problem in 20 min..that's why I wanted someone to teach me real quick..u know what i mean?

OpenStudy (asnaseer):

I know it can be frustrating at the start, but you put in some time to practice first then the rewards will be big for you. :)

OpenStudy (anonymous):

yeah you are right

OpenStudy (asnaseer):

ok, because you are cooperating here, I will do you a favour and show you how this particular one is done in the hope that you learn something :)

OpenStudy (anonymous):

ok thank you very much! :)

OpenStudy (asnaseer):

so, we had these equations to solve: t+u=9 10t+u+u=24 the first step was to look at just the first equation which is: t + u = 9 now, lets subtract u from both sides: t + u - u = 9 - u which leads to: t = 9 - u understand so far?

OpenStudy (anonymous):

ok, yes i do

OpenStudy (asnaseer):

good, so now we just need to substitute this expression for t into the second equation which was: 10t + u + u = 74 to get: 10(9 - u) + u + u = 74 expanding this we get: 90 - 10u + u + u = 74 then collect the terms involving u to get: 90 - 8u = 74 now lets add 8u to both sides to get: 90 - 8u + 8u = 74 + 8u which simplifies to: 90 = 74 + 8u now subtract 74 from both sides to get: 90 - 74 = 74 + 8u - 74 which simplifies to: 16 = 8u now divide both sides by 8 to get: 2 = u

OpenStudy (asnaseer):

BTW: I have corrected the second equation. it should not be: 10t+u+u=24 but this: 10t+u+u=74

OpenStudy (anonymous):

ok i see

OpenStudy (asnaseer):

so now we know u=2 we can use the expression we had for t: t = 9 - u = 9 - 2 = 7 so your two digit number was 72

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

and you can check this for correctness, the sum of the digits is = 7 + 2 = 9 number plus units digit = 72 + 2 = 74 which matches what the question stated

OpenStudy (asnaseer):

hope that helps. :)

OpenStudy (anonymous):

alright thank you so much!!! :)

OpenStudy (asnaseer):

yw :)